OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) a. 2x+3y=1 b. 2x-3y=7 c. 3x+27=4 d. 3x-2y=7.

OpenStudy (koikkara):

\(\color{darkgreen}{~\huge\star~``Welcome~to~Open~Study~!!"~\star~}\)

OpenStudy (anonymous):

let the equation be 2x+3y=m (m is the parameter) 2(9)+3(-3)-m=0 18-9-m=0 m=9

OpenStudy (anonymous):

thanks cody_123